OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form:

OpenStudy (anonymous):

Worst laptop in the world..

OpenStudy (anonymous):

\[\sqrt{3x} + 3\sqrt{3x} - 2\sqrt{3x} - 10\]

OpenStudy (anonymous):

I was thinking: \[2\sqrt{3x} - 10\]

OpenStudy (ashley_f97):

That's what I got

OpenStudy (anonymous):

Damn, I'm smart. Lol thanks.

OpenStudy (ashley_f97):

No problem

OpenStudy (anonymous):

;D bye.